the sides of square are 10 inches long. if you increase each side by 50% how many square inches will the area of the square grow.
Found 2 solutions by ewatrrr, lwsshak3:
Answer by ewatrrr(24785)

A = (10in)(10in) = 100in^2
...
increasing each side by 50%
A = (15in)(15in) = 15^2 = 225in^2
...
growth is 125in^2 more

Answer by lwsshak3(11628)

the sides of square are 10 inches long. if you increase each side by 50% how many square inches will the area of the square grow.
***
Area of a square varies as the square of its sides.
Area of original square=10*10=100 sq in
Area of new square=(1.5)^2*100=2.25*100=225 sq in